Sol Summary – December 28th

Sol 10

Long EVA

David Murray

Mission Status: Everything is going well

Sol Activity Summary: We woke up at 0800, had breakfast, and switched off the generator. Today, the EVA crew got a chance to gain experience on the ATV’s and to explore Lith Canyon. Then, they completed the navigation exercise and returned to the hab. After lunch, I planted Moringa oleifera seeds in the greenhab and watered the plants, Tibo and Ilaria participated in John’s nerve block experience, and Arno checked our state of charge. The crew plans to eat dinner and finish the rest of the movie that we started last night before going to bed.

Look Ahead Plan: Record our answers to questions from children around the world

Anomalies in work: None

Weather: Sunny, warm

Crew Physical Status: Healthy

EVA: Exploration, sample collection

Reports to be filed: Journalist Report, Greenhab Report, EVA Request, Sol Summary, EVA report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site